### v3

@@ 4.2 Propositions and Lemmas of $\sigma_{i}$ neural networks
~$\widetilde{\mathcal{R}}\leq 2^{d+7}C_{10}(n,d)\|f\|_{W^{n,p}([a,b]^{d})}N^{-2(n-2)/d}L^{-2(n-2)/d}.$
Overall, we have shown that DSRNs can approximate functions in Sobolev spaces with the $W^{2,p}$ norm. Our method extends directly to $W^{m,p}$ for any integer $m\geq 1$. The key modification in the proof is the construction of a $C^{m}$ partition of unity $\{s_{\boldsymbol{m}}(\boldsymbol{x})\}_{\boldsymbol{m}\in\{1,2\}^{d}}$, which we achieve by designing each $s_{\boldsymbol{m}}(\boldsymbol{x})$ as a piecewise polynomial of sufficiently high degree. In particular, to build a $C^{3}$ partition of unity, one may take the one-dimensional “bump” function
[…]
@@
$s(x)\;=\;\sum_{k=0}^{5}(-1)^{k}\binom{5}{k}\,\Bigl{[}\tfrac{x}{3}+\tfrac{2}{3}-k\Bigr{]}_{+}^{5}.$
By Lemma 4.3, each $\mathrm{ReLU}^{5}$ can itself be realized by a $\sigma_{2}$-NN. Consequently, we obtain a DSRN that approximates the target function in the $W^{3,p}$ norm. Repeating this construction inductively yields Corollary 4.23.
Corollary 4.23.
For any $f\in W^{n,p}([a,b]^{d})$, $m\in\mathbb{N}$ with $m\geq 2$ and $1\leq p\leq+\infty$, any $N,L\in\mathbb{N}_{+}$ with $N\log_{2}L+2^{\left\lfloor\log_{2}N\right\rfloor}\geq\max\{d,n\}$ and $L\geq N,m<n$, there is a DSRN $\gamma(\boldsymbol{x})$ in $\mathcal{N}_{\eta_{1},\eta_{2}L\log_{2}L}$ with the width $\eta_{3}N\log_{2}N$ such that
$\|f(\boldsymbol{x})-\gamma(\boldsymbol{x})\|_{W^{m,p}([a,b]^{d})}\leq C_{11}(n,d)\|f\|_{W^{n,p}([a,b]^{d})}N^{-2(n-m)/d}L^{-2(n-m)/d},$
@@
Theorem 5.1.
Given any $\rho,C_{1},C_{2},C_{3},C_{4},J_{0}>0$ and $n,d\in\mathbb{N}^{+}$, there exist $N,L\in\mathbb{N}$ with $NL\geq J_{0}$ and $f$, satisfying for any $\phi\in\mathcal{N}_{C_{4},C_{2}L\log L}$ with the width smaller than $C_{1}N\log N$, we have
(47) $|\phi-f|_{W^{2,p}([a,b]^{d})}>C_{3}\|f\|_{W^{n,p}([a,b]^{d})}L^{-2(n-2)/d-\rho}N^{-2(n-2)/d-\rho}.$
In order to prove the optimality of Theorem 4.1, we need to bound the VC-dimension of second order derivative of deep neural networks (DNNs).
Definition 5.2 (VC-dimension [1]).
@@
The Vapnik–Chervonenkis dimension (VC-dimension) of $H$, denoted by $\text{VCdim}(H)$, is the largest $m$ such that $\Pi_{H}(m)=2^{m}$. For a class $\mathcal{G}$ of real-valued functions, define $\text{VCdim}(\mathcal{G}):=\text{VCdim}(\mathrm{sgn}(\mathcal{G}))$, where $\mathrm{sgn}(\mathcal{G}):=\{\mathrm{sgn}(f):f\in\mathcal{G}\}$ and $\mathrm{sgn}(x)=1[x>0]$.
Theorem 5.3.
@@
(49) $D^{2}\Phi:=\left\{\lambda(\boldsymbol{x})=D^{\boldsymbol{\alpha}}\phi:\phi\in\Phi,~{}|\boldsymbol{\alpha}|=2\right\},$
@@
$D^{\boldsymbol{\alpha}}=D^{\alpha_{1}}_{1}D^{\alpha_{2}}_{2}\ldots D^{\alpha_{d}}_{d}$
as the partial derivative where $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{d}]^{T}$ and $D_{i}$ is the weak derivative in the $i$-th variable.
Prior to proving Theorem 5.3, we introduce two lemmas.
Lemma 5.4 ([3],[2]).
@@
$P:=\left|\{\left(\mathrm{sgn}(P_{1}(\boldsymbol{a})),\ldots,\mathrm{sgn}(P_{M}(\boldsymbol{a}))\right):\boldsymbol{a}\in\mathbb{R}^{W}\}\right|,$
@@
$\phi=\boldsymbol{W}_{L+1}\sigma(\boldsymbol{W}_{L}\sigma(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)+\boldsymbol{b}_{L})+b_{L+1},$
@@
$\psi(\boldsymbol{x})=D_{i}\phi(\boldsymbol{x})=$ $\boldsymbol{W}_{L+1}\sigma^{\prime}(\boldsymbol{W}_{L}\sigma(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)+\boldsymbol{b}_{L})$
(50) $\cdot\boldsymbol{W}_{L}\sigma^{\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\ldots\boldsymbol{W}_{2}\sigma^{\prime}(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})(\boldsymbol{W}_{1})_{i},$
where $\boldsymbol{W}_{i}\in\mathbb{R}^{N_{i}\times N_{i-1}}$ ($(\boldsymbol{W})_{i}$ is $i$-th column of $\boldsymbol{W}$) and $\boldsymbol{b}_{i}\in\mathbb{R}^{N_{i}}$ are the weight matrix and the bias vector in the $i$-th linear transform in $\phi$, and $\sigma^{\prime}(\boldsymbol{x})=\operatorname{diag}(\sigma^{\prime}(x_{i}))$. Then we have
(51) $\lambda(\boldsymbol{x})=D_{j}D_{i}\phi(\boldsymbol{x})=\sum_{a=1}^{L+1}\lambda_{a}(\boldsymbol{x}).$
@@
$\lambda_{a}(\boldsymbol{x})=$ $\boldsymbol{W}_{L+1}\sigma^{\prime}(\boldsymbol{W}_{L}\sigma(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)+\boldsymbol{b}_{L})\cdot\boldsymbol{W}_{L}\sigma^{\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\ldots$
$\cdot[\boldsymbol{W}_{a}\sigma^{\prime\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\cdot\boldsymbol{W}_{a-1}\sigma^{\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\boldsymbol{W}_{2}\sigma^{\prime}(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})(\boldsymbol{W}_{1})_{j}]$
(52) $\cdot\ldots\boldsymbol{W}_{a-1}\sigma^{\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\boldsymbol{W}_{2}\sigma^{\prime}(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})(\boldsymbol{W}_{1})_{i},$
where $\sigma^{\prime\prime}(\boldsymbol{x})=\operatorname{diag}(\sigma^{\prime\prime}(x_{i}))$ is a three-order tensor. Denote $W_{i}$ as the number of parameters in $\boldsymbol{W}_{i},\boldsymbol{b}_{i}$, i.e., $W_{i}=N_{i}N_{i-1}+N_{i}$.
Figure 5: The structure of $\lambda_{a}(\boldsymbol{x})$.
Note that $\sigma$ can either be the ReLU or the ReLU squared function. If $\sigma(x)$ is the ReLU function, then its second derivative, $\sigma^{\prime\prime}(x)$, will be the delta function, which is defined as $\sigma^{\prime\prime}(x)=0$ for $x\neq 0$ and $\sigma^{\prime\prime}(0)=\infty$, with $\int_{\mathbb{R}}\sigma^{\prime\prime}(x)\,\mathrm{d}x=1$. This delta function serves as the generalization function. However, since we assume that $\phi\in W^{2}([a,b]^{d})$, any delta functions must either disappear or be omitted in the $\lambda_{a}(\boldsymbol{x})$ equation, as shown in Eq. (52).
Let $\boldsymbol{x}\in\mathbb{R}^{d}$ be an input and $\boldsymbol{\theta}\in\mathbb{R}^{W}$ be a parameter vector in $\psi$. We denote the output of $\psi$ with input $\boldsymbol{x}$ and parameter vector $\boldsymbol{\theta}$ as $f(\boldsymbol{x},\boldsymbol{\theta})$. For fixed $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{m}$ in $\mathbb{R}^{d}$, we aim to bound
(53) $K:=\left|\{\left(\mathrm{sgn}(f(\boldsymbol{x}_{1},\boldsymbol{\theta})),\ldots,\mathrm{sgn}(f(\boldsymbol{x}_{m},\boldsymbol{\theta}))\right):\boldsymbol{\theta}\in\mathbb{R}^{W}\}\right|.$
@@
$K\leq\sum_{i=1}^{T}\left|\{\left(\mathrm{sgn}(f(\boldsymbol{x}_{1},\boldsymbol{\theta})),\ldots,\mathrm{sgn}(f(\boldsymbol{x}_{m},\boldsymbol{\theta}))\right):\boldsymbol{\theta}\in P_{i}\}\right|.$
We choose the partition such that within each region $P_{i}$, the functions $f(\boldsymbol{x}_{j},\cdot)$ are all fixed polynomials of bounded degree. This allows us to bound each term in the sum using Lemma 5.4.
We define a sequence of sets of functions $\{\mathbb{F}_{j}\}_{j=0}^{L}$ with respect to parameters $\boldsymbol{\theta}\in\mathbb{R}^{W}$:
[…]
$\boldsymbol{W}_{3}\sigma(\boldsymbol{W}_{2}\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})+\boldsymbol{b}_{2})+\boldsymbol{b}_{3}\}\cup\mathbb{F}_{1}$
@@
[…]
(54) $\boldsymbol{W}_{L+1}\sigma^{\prime\prime}(\boldsymbol{W}_{L}\sigma_{1}(\ldots\sigma_{1}(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)+\boldsymbol{b}_{L})\}\cup\mathbb{F}_{L-1}.$
@@
2. For each $n=0,\ldots,L^{*}$, each element $S$ of $\mathcal{S}_{n}$, when $\boldsymbol{\theta}$ varies in $S$, the output of each term in $\mathbb{F}_{n}$ is a fixed polynomial function in $\sum_{i=1}^{n}W_{i}$ variables of $\boldsymbol{\theta}$, with a total degree no more than $1+n2^{\max\{0,n-L^{*}\}}$.
We define $\mathcal{S}_{0}=\{\mathbb{R}^{W}\}$, which satisfies properties 1,2 above, since $\boldsymbol{W}_{1}\boldsymbol{x}_{j}+\boldsymbol{b}_{1}$ and $(\boldsymbol{W}_{1})_{i}$ for all $i=1,\ldots,d$ are affine functions of $\boldsymbol{W}_{1},\boldsymbol{b}_{1}$.
For each $n=0,\ldots,L$, to define $\mathcal{S}_{n}$, we use the last term of $\mathbb{F}_{n-1}$ as inputs for the new terms in $\mathbb{F}_{n}$. Assuming that $\mathcal{S}_{0},\mathcal{S}_{1},\ldots,\mathcal{S}_{n-1}$ have already been defined, we observe that the last two or three terms are new additions to $\mathbb{F}_{n}$ when comparing it to $\mathbb{F}_{n-1}$. Therefore, all elements in $\mathbb{F}_{n}$ except the $\mathbb{F}_{n}\backslash\mathbb{F}_{n-1}$ are fixed polynomial functions in $W_{n}$ variables of $\boldsymbol{\theta}$, with a total degree no greater than $1+(n-1)2^{\max\{0,n-1-L^{*}\}}$ when $\boldsymbol{\theta}$ varies in $S\in\mathcal{S}_{n}$. This is because $\mathcal{S}_{n}$ is a finer partition than $\mathcal{S}_{n-1}$.
We denote $p_{\boldsymbol{x}_{j},n-1,S,k}(\boldsymbol{\theta})$ as the output of the $k$-th node in the last term of $\mathbb{F}_{n-1}$ in response to $\boldsymbol{x}_{j}$ when $\boldsymbol{\theta}\in S$. The collection of polynomials
$\{p_{\boldsymbol{x}_{j},n-1,S,k}(\boldsymbol{\theta}):j=1,\ldots,m,~{}k=1,\ldots,N_{n}\}$
can attain at most $2\left(\frac{2em(1+(n-1)2^{\max\{0,n-1-L^{*}\}})N_{n}}{\sum_{i=1}^{n}W_{i}}\right)^{\sum_{i=1}^{n}W_{i}}$ distinct sign patterns when $\boldsymbol{\theta}\in S$ due to Lemma 5.4 for sufficiently large $m$. Therefore, we can divide $S$ into
@@
parts, each having the property that $p_{\boldsymbol{x}_{j},n-1,S,k}(\boldsymbol{\theta})$ does not change sign within the subregion. By performing this for all $S\in\mathcal{S}_{n-1}$, we obtain the desired partition $\mathcal{S}_{n}$. This division ensures that the required property 1 is satisfied.
Additionally, since the input to the last two terms in $\mathbb{F}_{n}$ is $p_{\boldsymbol{x}_{j},n-1,S,k}(\boldsymbol{\theta})$, and we have shown that the sign of this input will not change in each region of $\mathcal{S}_{n}$, it follows that the output of the last two terms in $\mathbb{F}_{n}$ is also a polynomial without breakpoints in each element of $\mathcal{S}_{n}$, therefore, the required property 2 is satisfied.
Due to the structure of $\lambda$, it is a polynomial function in $\sum_{i=1}^{L+1}W_{i}$ variables of $\boldsymbol{\theta}\in S\in\mathcal{S}_{L}$, of total degree no more than
@@
|$\left|\{\left(\mathrm{sgn}(f(\boldsymbol{x}_{1},\boldsymbol{\theta})),\ldots,\mathrm{sgn}(f(\boldsymbol{x}_{m},\boldsymbol{\theta}))\right):\boldsymbol{\theta}\in S\}\right|\leq 2\left(2emd_{2}/\sum_{i=1}^{L+1}W_{i}\right)^{\sum_{i=1}^{L+1}W_{i}}.$
@@
(55) $\leq$ $2^{L+3}\left(\frac{4em(L^{C+3})N}{U}\right)^{U},$
@@
(56) $2^{\text{VCdim}(D^{2}\Phi)}\leq 2^{L+3}\left(\frac{4e\text{VCdim}(D^{2}\Phi)L^{C+3}N}{U}\right)^{U}.$
Due to Lemma 5.5, we obtain that
(57) $\text{VCdim}(D^{2}\Phi)=\mathcal{O}(N^{2}L^{2}\log_{2}L\log_{2}N)$
since $U=\mathcal{O}(N^{2}L^{2})$.

### v4

@@ 4.3.1 Proof of Theorem 2
$\widetilde{\mathcal{R}}\leq 2^{d+7}C_{3}(n,d,a,b)\|f\|_{W^{n,p}([a,b]^{d})}N^{-2(n-2)/d}L^{-2(n-2)/d}.$
Overall, we have shown that DSRNs can approximate functions in Sobolev spaces with the $W^{2,p}$ norm. Our method extends directly to $W^{m,p}$ for any integer $m\geq 1$. The key modification in the proof is the construction of a $C^{m}$ partition of unity $\{s_{\bm{m}}(\bm{x})\}_{\bm{m}\in\{1,2\}^{d}}$, which we achieve by designing each $s_{\bm{m}}(\bm{x})$ as a piecewise polynomial of sufficiently high degree. In particular, to build a $C^{3}$ partition of unity, one may take the one-dimensional “bump” function
$s(x)\;=\;\begin{cases}0,&x\leq 0,\\ 6x^{5}\;-\;15x^{4}\;+\;10x^{3},&0<x<1,\\ 1,&1\leq x\leq 2,\\ 6(3-x)^{5}\;-\;15(3-x)^{4}\;+\;10(3-x)^{3},&2<x<3,\\ 0,&x\geq 3.\end{cases}$
@@
$s(x)\;=\;\sum_{k=0}^{5}(-1)^{k}\binom{5}{k}\,\operatorname{ReLU}^{5}\Bigl{(}\tfrac{x}{3}+\tfrac{2}{3}-k\Bigr{)}.$
By Lemma 3, each $\mathrm{ReLU}^{5}$ can itself be realized by a $\sigma_{2}$-NN. Consequently, we obtain a DSRN that approximates the target function in the $W^{3,p}$ norm. Repeating this construction inductively yields Corollary 8.
Corollary 8.
For any $f\in W^{n,p}([a,b]^{d})$, $m\in\mathbb{N}$ with $m\geq 2$ and $1\leq p\leq+\infty$, any $N,L\in\mathbb{N}_{+}$ with $N\log_{2}L+2^{\left\lfloor\log_{2}N\right\rfloor}\geq\max\{d,n\}$ and $L\geq N,m<n$, there is a DSRN $\gamma(\bm{x})$ in $\mathcal{N}_{\eta_{1},\eta_{2}L\log_{2}L}$ with the width $\eta_{3}N\log_{2}N$ such that
$\|f(\bm{x})-\gamma(\bm{x})\|_{W^{m,p}([a,b]^{d})}\leq C_{4}(n,d,a,b)\|f\|_{W^{n,p}([a,b]^{d})}N^{-2(n-m)/d}L^{-2(n-m)/d},$
@@ 5 Optimality of Theorem 2
Theorem 9.
Given any $\rho,D_{1},D_{2},D_{3},D_{4},J_{0}>0$ and $n,d\in\mathbb{N}^{+}$, there exist $N,L\in\mathbb{N}$ with $NL\geq J_{0}$ and $f$, satisfying for any $\phi\in\mathcal{N}_{D_{1},D_{2}L\log L}$ with the width smaller than $D_{3}N\log N$, we have
$|\phi-f|_{W^{2,p}([a,b]^{d})}>D_{4}\|f\|_{W^{n,p}([a,b]^{d})}L^{-2(n-2)/d-\rho}N^{-2(n-2)/d-\rho}.$ (5.1)
In order to prove the optimality of Theorem 2, we need to bound the VC-dimension of second order derivative of deep neural networks (DNNs).
Definition 6 (VC-dimension [1]).
@@
The Vapnik–Chervonenkis dimension (VC-dimension) of $H$, denoted by $\operatorname{VCdim}(H)$, is the largest $m$ such that $\Pi_{H}(m)=2^{m}$. For a class $\mathcal{G}$ of real-valued functions, define $\operatorname{VCdim}(\mathcal{G}):=\operatorname{VCdim}(\mathrm{sgn}(\mathcal{G}))$, where $\mathrm{sgn}(\mathcal{G}):=\{\mathrm{sgn}(f):f\in\mathcal{G}\}$ and $\mathrm{sgn}(x)=1[x>0]$.
Theorem 10.
@@
$D^{2}\Phi:=\left\{\lambda(\bm{x})=D^{\bm{\alpha}}\phi:\phi\in\Phi,~|\bm{\alpha}|=2\right\},$ (5.3)
@@
$D^{\bm{\alpha}}=D^{\alpha_{1}}_{1}D^{\alpha_{2}}_{2}\ldots D^{\alpha_{d}}_{d}$
as the partial derivative where $\bm{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{d}]^{T}$ and $D_{i}$ is the weak derivative in the $i$-th variable.
Prior to proving Theorem 10, we introduce two lemmas.
Lemma 11 ([3],[2]).
@@
$P:=\left|\{\left(\mathrm{sgn}(P_{1}(\bm{a})),\ldots,\mathrm{sgn}(P_{M}(\bm{a}))\right):\bm{a}\in\mathbb{R}^{W}\}\right|,$
@@
$\phi=\bm{W}_{L+1}\sigma(\bm{W}_{L}\sigma(\ldots\sigma(\bm{W}_{1}\bm{x}+\bm{b}_{1})\ldots)+\bm{b}_{L})+b_{L+1},$
@@
$\psi(\bm{x})=D_{i}\phi(\bm{x})=$ $\bm{W}_{L+1}\sigma^{\prime}(\bm{W}_{L}\sigma(\ldots\sigma(\bm{W}_{1}\bm{x}+\bm{b}_{1})\ldots)+\bm{b}_{L})$
$\cdot\bm{W}_{L}\sigma^{\prime}(\ldots\sigma(\bm{W}_{1}\bm{x}+\bm{b}_{1})\ldots)\ldots\bm{W}_{2}\sigma^{\prime}(\bm{W}_{1}\bm{x}+\bm{b}_{1})(\bm{W}_{1})_{i},$ (5.4)
where $\bm{W}_{i}\in\mathbb{R}^{N_{i}\times N_{i-1}}$ ($(\bm{W})_{i}$ is $i$-th column of $\bm{W}$) and $\bm{b}_{i}\in\mathbb{R}^{N_{i}}$ are the weight matrix and the bias vector in the $i$-th linear transform in $\phi$, and $\sigma^{\prime}(\bm{x})=\mathrm{diag}(\sigma^{\prime}(x_{i}))$. Then we have
$\lambda(\bm{x})=D_{j}D_{i}\phi(\bm{x})=\sum_{a=1}^{L+1}\lambda_{a}(\bm{x}).$
@@
$\lambda_{a}(\bm{x})=\bm{W}_{L+1}\sigma^{\prime}(\bm{W}_{L}\sigma(\ldots\sigma(\bm{W}_{1}\bm{x}+\bm{b}_{1})\ldots)+\bm{b}_{L})\cdot\bm{W}_{L}\sigma^{\prime}(\ldots\sigma(\bm{W}_{1}\bm{x}+\bm{b}_{1})\ldots)\ldots$
$\cdot[\bm{W}_{a}\sigma^{\prime\prime}(\ldots\sigma(\bm{W}_{1}\bm{x}+\bm{b}_{1})\ldots)\cdot\bm{W}_{a-1}\sigma^{\prime}(\ldots\sigma(\bm{W}_{1}\bm{x}+\bm{b}_{1})\ldots)\bm{W}_{2}\sigma^{\prime}(\bm{W}_{1}\bm{x}+\bm{b}_{1})(\bm{W}_{1})_{j}]$
$\cdot\ldots\bm{W}_{a-1}\sigma^{\prime}(\ldots\sigma(\bm{W}_{1}\bm{x}+\bm{b}_{1})\ldots)\bm{W}_{2}\sigma^{\prime}(\bm{W}_{1}\bm{x}+\bm{b}_{1})(\bm{W}_{1})_{i},$
where $\sigma^{\prime\prime}(\bm{x})=\mathrm{diag}(\sigma^{\prime\prime}(x_{i}))$ is a three-order tensor. The multiplication at here is defined as follows: Let a third-order tensor $\mathcal{T}\in\mathbb{R}^{I\times J\times K}$ and three matrices $A\in\mathbb{R}^{P\times I}$, $B\in\mathbb{R}^{Q\times J}$, and $C\in\mathbb{R}^{R\times K}$. The multiplication of the tensor with these matrices is defined as
$\mathcal{T}^{\prime}=\mathcal{T}\times_{1}A\times_{2}B\times_{3}C:=(\mathcal{T}^{\prime}_{pqr})_{p,q,r}\;\;\in\mathbb{R}^{P\times Q\times R},$
where
$\mathcal{T}^{\prime}_{pqr}=\sum_{i=1}^{I}\sum_{j=1}^{J}\sum_{k=1}^{K}A_{pi}\,B_{qj}\,C_{rk}\,\mathcal{T}_{ijk}.$
This operation applies a linear transformation to each mode of the tensor $\mathcal{T}$.
Figure 6: The structure of $\lambda_{a}(\bm{x})$.
Note that $\sigma$ can either be the ReLU or the ReLU squared function. If $\sigma(x)$ is the ReLU function, then its second derivative, $\sigma^{\prime\prime}(x)$, will be the delta distribution, which is defined as $\sigma^{\prime\prime}(x)=0$ for $x\neq 0$ and $\sigma^{\prime\prime}(0)=\infty$, with $\int_{\mathbb{R}}\sigma^{\prime\prime}(x)\,\mathrm{d}x=1$. This delta function serves as the generalization function (a distribution. However, since we assume that $\phi\in W^{2}([a,b]^{d})$, any delta functions must either disappear or be omitted in the $\lambda_{a}(\bm{x})$ equation.
Denote $W_{i}$ as the number of parameters in $\bm{W}_{i},\bm{b}_{i}$, i.e., $W_{i}=N_{i}N_{i-1}+N_{i}$. Let $\bm{x}\in\mathbb{R}^{d}$ be an input and $\bm{\theta}\in\mathbb{R}^{W}$ be a parameter vector in $\psi$. We denote the output of $\psi$ with input $\bm{x}$ and parameter vector $\bm{\theta}$ as $f(\bm{x},\bm{\theta})$. For fixed $\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{m}$ in $\mathbb{R}^{d}$, we aim to bound
$K:=\left|\{\left(\mathrm{sgn}(f(\bm{x}_{1},\bm{\theta})),\ldots,\mathrm{sgn}(f(\bm{x}_{m},\bm{\theta}))\right):\bm{\theta}\in\mathbb{R}^{W}\}\right|.$ (5.5)
@@
$K\leq\sum_{i=1}^{T}\left|\{\left(\mathrm{sgn}(f(\bm{x}_{1},\bm{\theta})),\ldots,\mathrm{sgn}(f(\bm{x}_{m},\bm{\theta}))\right):\bm{\theta}\in P_{i}\}\right|.$
We choose the partition such that within each region $P_{i}$, the functions $f(\bm{x}_{j},\cdot)$ are all fixed polynomials of bounded degree. This allows us to bound each term in the sum using Lemma 11.
We define a sequence of sets of functions $\{\mathbb{F}_{j}\}_{j=0}^{L}$ with respect to parameters $\bm{\theta}\in\mathbb{R}^{W}$:
$\mathbb{F}_{0}:=\cup_{j=1}^{m}\{$ $(\bm{W}_{1})_{1},(\bm{W}_{1})_{2},\ldots,(\bm{W}_{1})_{d},\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1}\}$
$\mathbb{F}_{1}:=\cup_{j=1}^{m}\{$ $\bm{W}_{2}\sigma^{\prime\prime}(\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1}),\bm{W}_{2}\sigma^{\prime}(\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1}),\bm{W}_{2}\sigma(\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1})+\bm{b}_{2}\}\cup\mathbb{F}_{0}$
$\mathbb{F}_{2}:=\cup_{j=1}^{m}\{$ $\bm{W}_{3}\sigma^{\prime\prime}(\bm{W}_{2}\sigma(\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1})+\bm{b}_{2}),\bm{W}_{3}\sigma^{\prime}(\bm{W}_{2}\sigma(\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1})+\bm{b}_{2}),$
$\bm{W}_{3}\sigma(\bm{W}_{2}\sigma(\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1})+\bm{b}_{2})+\bm{b}_{3}\}\cup\mathbb{F}_{1}$
@@
$\mathbb{F}_{L}:=\cup_{j=1}^{m}\{$ $\bm{W}_{L+1}\sigma^{\prime}(\bm{W}_{L}\sigma_{1}(\ldots\sigma_{1}(\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1})\ldots)+\bm{b}_{L}),$
$\bm{W}_{L+1}\sigma^{\prime\prime}(\bm{W}_{L}\sigma_{1}(\ldots\sigma_{1}(\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1})\ldots)+\bm{b}_{L})\}\cup\mathbb{F}_{L-1}.$
@@
2. For each $n=0,\ldots,L^{*}$, each element $S$ of $\mathcal{S}_{n}$, when $\bm{\theta}$ varies in $S$, the output of each term in $\mathbb{F}_{n}$ is a fixed polynomial function in $\sum_{i=1}^{n}W_{i}$ variables of $\bm{\theta}$, with a total degree no more than $1+n2^{\max\{0,n-L^{*}\}}$.
We define $\mathcal{S}_{0}=\{\mathbb{R}^{W}\}$, which satisfies properties 1,2 above, since $\bm{W}_{1}\bm{x}_{j}+\bm{b}_{1}$ and $(\bm{W}_{1})_{i}$ for all $i=1,\ldots,d$ are affine functions of $\bm{W}_{1},\bm{b}_{1}$.
For each $n=0,\ldots,L$, to define $\mathcal{S}_{n}$, we use the last term of $\mathbb{F}_{n-1}$ as inputs for the new terms in $\mathbb{F}_{n}$. All elements in $\mathbb{F}_{n}$ except the $\mathbb{F}_{n}\backslash\mathbb{F}_{n-1}$ are fixed polynomial functions in $W_{n}$ variables of $\bm{\theta}$, with a total degree no greater than $1+(n-1)2^{\max\{0,n-1-L^{*}\}}$ when $\bm{\theta}$ varies in $S\in\mathcal{S}_{n}$. This is because $\mathcal{S}_{n}$ is a finer partition than $\mathcal{S}_{n-1}$.
We denote $p_{\bm{x}_{j},n-1,S,k}(\bm{\theta})$ as the output of the $k$-th node in the last term of $\mathbb{F}_{n-1}$ in response to $\bm{x}_{j}$ when $\bm{\theta}\in S$. The collection of polynomials
$\{p_{\bm{x}_{j},n-1,S,k}(\bm{\theta}):j=1,\ldots,m,~k=1,\ldots,N_{n}\}$
can attain at most $2\left(\frac{2em(1+(n-1)2^{\max\{0,n-1-L^{*}\}})N_{n}}{\sum_{i=1}^{n}W_{i}}\right)^{\sum_{i=1}^{n}W_{i}}$ distinct sign patterns when $\bm{\theta}\in S$ due to Lemma 11 for sufficiently large $m$. Therefore, we can divide $S$ into
@@
parts, each having the property that $p_{\bm{x}_{j},n-1,S,k}(\bm{\theta})$ does not change sign within the subregion. By performing this for all $S\in\mathcal{S}_{n-1}$, we obtain the desired partition $\mathcal{S}_{n}$. This division ensures that the required property 1 is satisfied.
Additionally, since the input to terms in $\mathbb{F}_{n}\backslash\mathbb{F}_{n-1}$ is $p_{\bm{x}_{j},n-1,S,k}(\bm{\theta})$, and we have shown that the sign of this input will not change in each region of $\mathcal{S}_{n}$, it follows that the output of the terms in $\mathbb{F}_{n}\backslash\mathbb{F}_{n-1}$ is also a polynomial without breakpoints in each element of $\mathcal{S}_{n}$, therefore, the required property 2 is satisfied. Due to the structure of $\lambda$, it is a polynomial function in $\sum_{i=1}^{L+1}W_{i}$ variables of $\bm{\theta}\in S\in\mathcal{S}_{L}$, of total degree no more than
@@
$\left|\{\left(\mathrm{sgn}(f(\bm{x}_{1},\bm{\theta})),\ldots,\mathrm{sgn}(f(\bm{x}_{m},\bm{\theta}))\right):\bm{\theta}\in S\}\right|\leq 2\left(2emd_{2}/\sum_{i=1}^{L+1}W_{i}\right)^{\sum_{i=1}^{L+1}W_{i}}.$
@@
$\leq$ $2^{L+3}\left(\frac{4em(L^{C+3})N}{U}\right)^{U},$
@@
$2^{\operatorname{VCdim}(D^{2}\Phi)}\leq 2^{L+3}\left(\frac{4e\operatorname{VCdim}(D^{2}\Phi)L^{C+3}N}{U}\right)^{U}.$
Due to Lemma 12, we obtain that
$\operatorname{VCdim}(D^{2}\Phi)=\mathcal{O}(N^{2}L^{2}\log_{2}L\log_{2}N)$
since $U=\mathcal{O}(N^{2}L^{2})$. ∎